# Cooperative Bistatic ISAC Systems for Low-Altitude Economy

### v1

@@ V-C Performance for 3D Position and Velocity Estimation
Fig. 8 shows the 3D position and velocity estimation RMSE performance of both schemes under the scenario with $K=3$ targets. A scheme termed Single-pair is also introduced as the localization performance baseline. For any target, this scheme selects the basic position estimate generated by its nearest tBS and rBS as the optimal result. Therefore, the performance of Single-pair scheme represents the upper bound of localization performance for non-cooperative bistatic sensing schemes under the considered scenario. The results in Fig. 8(a) reveal that significant accuracy improvements are achieved through multi-BS cooperation at identical transmit power levels (50 dBm or 60 dBm). Even with the minimum configuration ($N{\mathrm{t}}\times N{\mathrm{r}}=2$ tBS-rBS pairs) and the Averaging scheme, cooperative localization outperforms its non-cooperative counterpart. In addition, it can be observed from Fig. 8 that the proposed scheme substantially enhances multi-target localization accuracy compared to the baselines, thereby enabling higher 3D velocity estimation accuracy. This performance gain becomes increasingly significant with the addition of cooperative BSs, highlighting the advantages of our proposed soft fusion algorithm over traditional hard fusion methods.

### v2

@@ V-C Performance for 3D Position and Velocity Estimation
Fig. 9 shows the 3D position and velocity estimation RMSE performance of both schemes under the scenario with $K=3$ targets. A scheme termed Single-pair is also introduced as the localization performance baseline. For any target, this scheme selects the basic position estimate generated by its nearest tBS and rBS as the optimal result. Therefore, its performance represents the upper bound of localization performance for non-cooperative bistatic sensing schemes under the considered scenario. The results in Fig. 9(a) reveal that significant accuracy improvements are achieved through multi-BS cooperation at identical transmit power levels (45 dBm or 55 dBm). Even with the minimum configuration ($N{\mathrm{t}}\times N{\mathrm{r}}=2$ tBS-rBS pairs) and the Averaging scheme, cooperative localization outperforms its non-cooperative counterpart. In addition, the results in Fig. 9 shows that the proposed scheme substantially enhances multi-target localization accuracy compared to the baselines, thereby enabling higher 3D velocity estimation accuracy. This performance gain becomes increasingly significant with the addition of cooperative BSs, highlighting the advantages of our proposed soft fusion algorithm over traditional hard fusion methods.